OpenStudy (anikate):

State the range of this function: y=2x-7/2

OpenStudy (anonymous):

the range is all reals

OpenStudy (anikate):

huh?

OpenStudy (anonymous):

(-infinity, 0) (0, infinity)

OpenStudy (anikate):

what?

OpenStudy (anikate):

i dont get it

OpenStudy (anikate):

keep goig i'l try to figure it out

OpenStudy (anonymous):

There are no restrictions to the range

OpenStudy (anikate):

um... need to find range is all i no

OpenStudy (anonymous):

Okay, well you could always substitute 0 for the x value and work out the whole question from there. However, the answer I gave you earlier should be correct

OpenStudy (anikate):

what was the answer u gave earlier?

OpenStudy (anikate):

-3.5?

OpenStudy (anonymous):

All real numbers. The domain would be all real numbers, and so should the range

OpenStudy (anikate):

what r real numbers?

OpenStudy (anonymous):

Or, you could set y= 0 and work out the whole equation

OpenStudy (anikate):

then what would be my answer?

OpenStudy (anikate):

-3.5?

OpenStudy (anonymous):

For instance, (0= 2x - 7/ 2 ) x 2 0= 4x-14 14=4x 14/4=x

OpenStudy (anonymous):

That's solving for x

OpenStudy (anikate):

range is always y right?

OpenStudy (anonymous):

Substitution with x and zero would look like this: y= 2(0) -7/2 y= -5/2

OpenStudy (anonymous):

that is the range when solving the equation

OpenStudy (anikate):

actually it would be -3.5

OpenStudy (anonymous):

Well, if you were trying to be precise, then yes. However, most math classes perfer that you leave it in the fractional form.

OpenStudy (anikate):

idk the range when soling equatyion. and i know

OpenStudy (anonymous):

also it should be -2.5, not -3.5

OpenStudy (anikate):

how ?

OpenStudy (anonymous):

2-7 = -5 -5 divided by 2 = -2.5

OpenStudy (anikate):

y=2(0)-7/2 y=0-7/2 y=-3.5

OpenStudy (anonymous):

oh, I just got served! You are absolutely right my dear. Thanks for catching my mistake ;}

OpenStudy (anikate):

oh it's k np

OpenStudy (anonymous):

So, do you have your answer?

OpenStudy (anikate):

-3.5?

OpenStudy (anonymous):

yes

OpenStudy (anikate):

yup thx :D

OpenStudy (anonymous):

No worries, glad I could help :)